Question 49
ProbabilityA bag contains 3 red and 2 white identical balls. If 2 balls are picked at random from the bag, one after the other and with replacement, find the probability that they are of different colours
- A.36/625
- B.16/625
- C.12/25
- D.13/25
Show the answer and explanation
Answer: C (12/25)
Explanation
P(red then white) + P(white then red) = 3/5 × 2/5 + 2/5 × 3/5 = 12/25.